a bullet with a 2.0 g mass 300 m/s velocity hits a wooden trolley with 1.5 kg mass which is the rest on a frictionless track. determine the final velocity of the wooden trolley.
1
Expert's answer
2020-09-28T08:05:16-0400

The bullet has the following momentum before the impact:


"p_b = 300m/s\cdot 0.002 kg = 0.6\space (m\cdot kg)/s"

Assume, that it transfers the whole momentum to the wooden trolley. Then the momentum of the trolley is:


"p_t = p_b = 1.5kg\cdot v_t"

where "v_t" is the final velocity of the trolley. Expressing "v_t" obtain:


"v_t = \dfrac{p_b}{1.5kg} = \dfrac{0.6\space (m\cdot kg)/s}{1.5kg} = 0.4m/s"

Answer. 0.4 m/s.